Can you help me find the intercept of -2x - y = 10
I have (10, 0), but am not sure if it should be reversed to (0, 10)
thank you, Ellen

Answer by HyperBrain(694)

That's a common error. The slope intercept of the line is
y=mx+b
m is the slope while
b is the y-coordinate of the y-intercept

Take note that the x-coordinate of the y-intercept is always zero. since, it is ON the y-axis.
Solve:
-2x-y=10
Substituting x=0,
-2(0)-y=10
-0-y=10
-y=10
y=-10

Oops! i think you got the wrong y-coordinate. The y-intercept is at (0,-10)

Power up,
HyperBrain!